# ESGenius: Benchmarking LLMs on Environmental, Social, and Governance (ESG) and Sustainability Knowledge

### v1

@@ Abstract
We introduce ESGenius, a comprehensive benchmark for evaluating and enhancing the proficiency of Large Language Models (LLMs) in Environmental, Social and Governance (ESG) and sustainability-focused question answering. ESGenius comprises two key components: (i) ESGenius-QA, a collection of 1,136 MCQs generated by LLMs and rigorously validated by domain experts, covering a broad range of ESG pillars and sustainability topics. Each question is systematically linked to its corresponding source text, enabling transparent evaluation and supporting Retrieval-Augmented Generation (RAG) methods; and (ii) ESGenius-Corpus, a meticulously curated repository of 231 foundational frameworks, standards, reports, and recommendation documents from 7 authoritative sources. Moreover, to fully assess the capabilities and adaptation potential of the model, we implement a rigorous two-stage evaluation protocol—Zero-Shot and RAG. Extensive experiments across 50 LLMs (ranging from 0.5B to 671B parameters) demonstrate that state-of-the-art models achieve only moderate performance in zero-shot settings, with accuracies typically around 55-70%, highlighting ESGenius’s challenging nature for LLMs in interdisciplinary contexts. However, models employing RAG demonstrate significant performance improvements, particularly for smaller models. For example,“DeepSeek-R1-Distill-Qwen 14B” improves from 63.82% (zero-shot) to 80.46% with RAG. These results demonstrate the necessity of grounding responses in authoritative sources for enhanced ESG understanding. To the best of our knowledge, ESGenius is the first benchmark curated for LLMs and the relevant enhancement technologies, focusing on ESG and sustainability topics.
@@
Chaoyue He1   Xin Zhou1   Yi Wu1   Xinjia Yu1   Yan Zhang1   Lei Zhang1 Di Wang1Shengfei Lyu1Hong Xu1Xiaoqiao Wang2Wei Liu2Chunyan Miao1 1Alibaba-NTU Global e-Sustainability CorpLab (ANGEL), Singapore, 2Alibaba Group, China GitHub: https://github.com/ANGEL-NTU/ESGenius Web Portal: https://angel-ntu.github.io/ESGenius
@@ 1 Introduction
Figure 1: Relationship between model size and zero-shot accuracy across 50 LLMs evaluated on the ESGenius benchmark. Model sizes are plotted on a log10 scale (in billions of parameters), with accuracy shown as percentages. There is a moderate positive correlation between model size and performance, suggesting larger models generally perform better. Dot colours denote 5 model families (DeepSeek, Gemma, GPT, Meta-Llama, Qwen). For proprietary API models, parameter counts are based on industry estimates (see Table 2 for details).
@@
Environmental, Social, and Governance (ESG) knowledge encompasses a vast domain of sustainability and corporate responsibility information that LLMs must effectively process to serve emerging business needs. This field spans across critical agendas from climate change and emissions tracking (IPCC, 2023), to workplace safety and human rights (Global Reporting Initiative, 2023). Such knowledge is codified in numerous technical frameworks and standards—major ones including GRI (Initiative, 2023), SASB (Board, 2023), TCFD (Board, 2017), ISSB(International Sustainability Standards Board, 2023), and CDP (CDP Worldwide, 2023)—which are constantly evolving.
@@
While LLMs show promise in processing complex ESG documents and providing relevant answers to user queries, their capabilities in this interdisciplinary domain remain largely unevaluated. Considering the high stakes involved, where incorrect responses about ESG requirements or metrics could lead to serious compliance violations or misguided sustainability initiatives, this lack of assessment is particularly problematic. However, there exists no comprehensive question-answering (QA) benchmark specifically designed to evaluate how well LLMs understand and reason about ESG concepts. Most existing QA datasets either completely omit ESG topics or address them superficially. This gap leaves researchers and practitioners without a reliable way to measure and improve LLMs’ ESG knowledge comprehension and question-answering abilities.
@@
To close this gap, we present ESGenius, a curated benchmark that targets MCQ-QA as the core evaluation task for ESG and sustainability knowledge. Our contributions are fourfold: (1) ESGenius Benchmark: A comprehensive evaluation framework comprising two integrated components: (i) ESGenius-QA, a collection of 1,136 MCQs across various ESG pillars and sustainability topics, generated through SOTA LLM approaches and validated by domain experts. Each question is explicitly mapped to supporting evidence from authoritative source texts, enabling transparent evaluation and facilitating RAG applications; and (ii) ESGenius-Corpus, a collection of 231 ESG documents and frameworks enabling efficient knowledge retrieval from 7 major authoritative sources. (2) Evaluation Protocol: We implement a comprehensive two-stage evaluation protocol consisting of zero-shot testing and RAG to systematically assess LLM capabilities. This assessment framework provides valuable insights into the current limitations and future potential of LLMs in understanding ESG.(3) Evaluation Analysis: Testing across a diverse set of 50 LLMs (ranging from 0.5B to 671B parameters) reveals significant performance gaps in zero-shot settings, with most models achieving 55-70% accuracy and the best model (o3) achieves a top score of 72.54%, as shown in Figure 1. However, models demonstrate substantial potential for improvement through RAG approaches, with “DeepSeek-R1-Distill-Qwen 14B” improving from 63.82% (zero-shot) to 80.46%. (4) Open Source Initiative: To foster community engagement and collaborative advancement, we have made our complete benchmark suite publicly available at https://github.com/ANGEL-NTU/ESGenius. This includes comprehensive documentation, evaluation code, model implementations, and the full ESGenius dataset. We also maintain an interactive web portal at https://angel-ntu.github.io/ESGenius/ featuring a real-time leaderboard and detailed performance visualizations through heatmaps (see Appendix Figures 7 and 8), enabling researchers to track progress and identify areas for improvement in ESG-focused language models.
@@
The remainder of this paper is organized as follows: Section 2 surveys prior efforts on ESG knowledge resources, question-answering benchmarks, and retrieval-augmented generation; Section 3 details the construction of the ESGenius benchmark, describing both the ESGenius-QA dataset and the ESGenius-Corpus; Section 4 outlines our experimental protocols—including zero-shot and RAG settings—and presents a comprehensive evaluation of 50 LLMs; Section 5 discusses the benchmark’s current limitations and avenues for future extension; and Section 6 concludes the paper and highlights directions for subsequent research.
@@ 2 Related Work
This section reviews relevant literature across three key areas in using LLMs for ESG benchmarking, including knowledge sources, QA benchmarks and RAG.
@@ 2.1 ESG & Sustainability Knowledge Resources
ESG data have traditionally been guided by voluntary reporting frameworks such as the Global Reporting Initiative (GRI) and the Sustainability Accounting Standards Board (SASB). These frameworks define taxonomies of ESG topics (e.g., emissions, labor practices, board governance) and indicators, but the unstructured text in corporate sustainability reports poses significant challenges for computational use. To impose structure, researchers have proposed ontologies and knowledge bases for ESG. For instance, Zhou and Perzylo (2023) introduced OntoSustain, aligning GRI and EU standards to capture key sustainability indicators, though broad coverage and automated population remain difficult.
Several ESG-focused text corpora have emerged to support NLP research in this domain. One recent example is SusGen-30K, a dataset of 30k instances spanning multiple financial NLP tasks (e.g., sentiment, Q&A) and an ESG report generation task Wu et al. (2024). Meanwhile, Chang et al. (2024) built a benchmark labeling sentences from corporate sustainability reports covering 36 ESG topics. These datasets remain relatively restricted, especially compared to general-domain corpora, and focus primarily on classification or generation tasks rather than MCQ question answering, thus motivating the need for broader coverage like ESGenius.
Prior NLP work on sustainability texts has largely centered around information extraction and text classification. Early methods relied on topic models or lexicon-based approaches Raghupathi and Raghupathi (2020), but these often fell short in capturing nuanced ESG terminology. More recent efforts leverage LLMs. For instance, Bronzini et al. (2023) employed a GPT-based pipeline to extract triples (company, action, ESG issue) from sustainability disclosures, populating a knowledge graph covering 500+ ESG topics. Similarly, Marodon et al. (2022) unveils SDG Prospector, an AI-driven DistilRoBERTa NLP tool that mines 237 public development banks’ 2016-2020 reports to expose which Sustainable Development Goals dominate their strategic narrative.
Researchers have also developed specialized language models for ESG. Webersinke et al. (2021) introduced ClimateBERT, continuing BERT’s pre-training on 1.6M climate-related paragraphs, enabling more effective classification of risk disclosures. In industry, Ni et al. (2023) proposed ChatReport, which uses GPT-based summarization and semantic search to assess a company’s report against TCFD recommendations. These domain-specific approaches illustrate the promise of LLMs for ESG tasks but also highlight the need for standardized benchmarks for a thorough evaluation.
@@ 2.2 QA & Evaluation Benchmarks
Question answering (QA) benchmarks have driven many NLP advancements over the past decade, starting with the TREC QA tracks Voorhees and Tice (2001) and continuing through datasets like SQuAD Rajpurkar et al. (2016). While these resources predominantly test general factual or reading comprehension, specialized QA datasets have emerged to measure higher-level reasoning and domain expertise. For example, OpenBookQA Mihaylov et al. (2018) challenged models to combine a small “open book” of scientific facts with common sense, while BoolQ Clark et al. (2019) featured yes/no queries requiring nuanced text understanding. Datasets such as DROP Dua et al. (2019) emphasize discrete reasoning skills, involving counting or arithmetic over paragraphs.
Further expansions include large-scale evaluation suites like MMLU (Massive Multitask Language Understanding) Hendrycks et al. (2021), covering dozens of subjects at college exam level, and domain-specific tasks such as PubMedQA Jin et al. (2019) for biomedical research. Despite these efforts, there remains a dearth of QA benchmarks devoted to ESG and/or sustainability. Thus, ESGenius fills in this gap with a large set of expert-validated MCQs spanning across environmental, social, and governance topics. By drawing from comprehensive frameworks (e.g., IPCC, GRI, SASB), ESGenius offers a high-stakes, real-world testbed that is specific to ESG.
@@ 2.3 RAG and Knowledge Grounding
When tackling specialized or dynamically updated domains, RAG has emerged as a powerful method for accurate and interpretable QA. Lewis et al. (2020) introduced RAG, an approach that combines a parametric language model with a non-parametric memory of documents. This approach improved performance on knowledge-intensive QA tasks (e.g., NaturalQuestions) relative to purely parametric models, as it allows the LLM to ground responses in retrieved evidence.
Other retrieval-generation frameworks such as Fusion-in-Decoder (Izacard and Grave, 2021) and REALM (Guu et al., 2020) illustrated diverse strategies for fetching relevant text chunks to guide the model’s reasoning. Recently, these methods have been applied to high-stakes sectors including finance and climate science. For instance, Vaghefi et al. (2023) integrated GPT-4 with a 3,000-page IPCC report to answer climate change questions more accurately. Similarly, Ni et al. (2023) used semantic search to retrieve TCFD-related segments from a company’s sustainability report, prompting an LLM to summarize potential climate risks. This grounding is critical in ESG contexts, where traceability and authoritative references are essential.
@@ 3 The ESGenius Benchmark
Developing reliable AI tools for the complex ESG and sustainability domain requires high-quality, domain-specific data. Our ESGenius benchmark provides a comprehensive foundation to evaluate and enhance LLMs in this context. It is designed to encompass the key dimensions of ESG, drawing from globally recognized standards and authoritative sources. The benchmark comprises two interconnected components: ESGenius-QA, a challenging question-answering dataset covering environmental, social, and governance topics with explicit linkages to textual evidence; and ESGenius-Corpus, encompassing a diverse collection of authoritative sources in ESG and sustainability. The detailed coverage of sources and statistics are presented in Appendix D.
@@ 3.1 ESGenius-QA
The ESGenius-QA dataset comprises 1,136 carefully curated multiple-choice questions (MCQs) with ground truth answers, validated by ESG domain experts. The questions systematically cover essential ESG topics across three core pillars: (i) Environmental: Climate change mitigation and adaptation, comprehensive carbon accounting (Scope 1, 2, 3 emissions), energy efficiency, water resource management, biodiversity conservation, waste reduction, and pollution control; (ii) Social: Labor practices and standards, human rights protection, diversity and inclusion initiatives, workplace health and safety, community engagement and impact assessment, data privacy and protection; (iii) Governance: Board structure and independence, executive compensation frameworks, business ethics and compliance, shareholder rights and engagement, enterprise risk management, and regulatory adherence. The dataset is generated through our specialized pipeline (detailed in Appendix A.1) and undergoes rigorous validation by domain experts to ensure accuracy and relevance, as described in 3.1.
@@
The ESGenius-QA dataset is structured to reflect the complexity and diversity of real-world ESG assessments through a standardized multiple-choice format. Each question contains four options (A-D) with a single correct answer, plus a dedicated “Not sure” option (Z) to assess model uncertainty. This design enables systematic evaluation of both performance and confidence. The dataset carefully balances conciseness with domain-specific precision: questions and options employ precise ESG terminology while maintaining clarity and focus, with contextual prompts provided where necessary for essential background. The specialized lexicon throughout authentically represents real-world ESG assessment scenarios. Table 1 presents a detailed token-level analysis of questions, options, and source text. Key terminology distributions are visualized through word clouds for the source text (Figure 4), questions (Figure 2), and answer options (Figure 3).
@@
Each question in ESGenius-QA is precisely mapped to relevant page(s) and text passages within authoritative source documents (GRI, SASB, IPCC, etc.) containing the information required for correct answers. ESG domain experts meticulously annotate these mappings, ensuring both accuracy and relevance. This explicit linking between questions and their supporting evidence in source texts enables effective RAG approaches.
@@
To ensure the highest quality of questions and answers, we implement a rigorous multi-stage validation process. After initial generation through our LLM pipeline mentioned in Appendix A.1, each question is assigned to domain experts. These experts carefully validate that: (1) the question-answer pair is correct according to authoritative sources, (2) the question is clearly worded and tests meaningful domain knowledge, (3) distractors are plausible but unambiguously incorrect, and (4) sufficient context is provided where needed. Questions failing any validation criteria are either refined or removed from the dataset.

### v2

@@ Abstract
We introduce ESGenius, a comprehensive benchmark for evaluating and enhancing the proficiency of Large Language Models (LLMs) in Environmental, Social, and Governance (ESG) and sustainability-focused question answering. ESGenius comprises two key components: (i) ESGenius-QA, a collection of 1,136 Multiple-Choice Questions (MCQs) generated by LLMs and rigorously validated by domain experts, covering a broad range of ESG pillars and sustainability topics. Each question is systematically linked to its corresponding source text, enabling transparent evaluation and supporting Retrieval-Augmented Generation (RAG) methods; and (ii) ESGenius-Corpus, a meticulously curated repository of 231 foundational frameworks, standards, reports, and recommendation documents from 7 authoritative sources. Moreover, to fully assess the capabilities and adaptation potential of LLMs, we implement a rigorous two-stage evaluation protocol—Zero-Shot and RAG. Extensive experiments across 50 LLMs (0.5B to 671B) demonstrate that state-of-the-art models achieve only moderate performance in zero-shot settings, with accuracies around 55–70%, highlighting a significant knowledge gap for LLMs in this specialized, interdisciplinary domain. However, models employing RAG demonstrate significant performance improvements, particularly for smaller models. For example, DeepSeek-R1-Distill-Qwen-14B improves from 63.82% (zero-shot) to 80.46% with RAG. These results demonstrate the necessity of grounding responses in authoritative sources for enhanced ESG understanding. To the best of our knowledge, ESGenius is the first comprehensive QA benchmark designed to rigorously evaluate LLMs on ESG and sustainability knowledge, providing a critical tool to advance trustworthy AI in this vital domain.
@@
Chaoyue He1   Xin Zhou1 thanks: Corresponding author  Yi Wu1   Xinjia Yu1   Yan Zhang1   Lei Zhang1   Di Wang1 Shengfei Lyu1Hong Xu1Xiaoqiao Wang2Wei Liu2Chunyan Miao1 1Alibaba-NTU Global e-Sustainability CorpLab (ANGEL), Singapore;   2Alibaba Group, China {cyhe, xin.zhou, xinjia.yu, zhang.yan, wangdi, shengfei.lyu, xuhong, ascymiao}@ntu.edu.sg {wuyi0614, leizhanzzl.1103}@gmail.com   {nerissa.wxq, weiliu.liuwei}@alibaba-inc.com GitHub: https://github.com/ANGEL-NTU/ESGenius Web Portal: https://angel-ntu.github.io/ESGenius
@@ 1 Introduction
Figure 2: Relationship between model size and zero-shot accuracy across 50 LLMs evaluated on the ESGenius benchmark. Model sizes are plotted on a $\log_{10}$ scale (in billions of parameters), with accuracy shown as percentages. There is a moderate positive correlation between model size and performance, suggesting larger models generally perform better. Dot colors denote five model families (DeepSeek, Gemma, GPT, Meta-Llama, Qwen). For proprietary API models, parameter counts are based on industry estimates (See Table 2 for details).
@@
Environmental, Social, and Governance (ESG) knowledge encompasses a vast domain of sustainability and corporate responsibility information that Large Language Models (LLMs) must effectively process to serve emerging business needs Zhou et al. (2024); Singh et al. (2025); Ong et al. (2025); Zou et al. (2025); Zhang et al. (2024); Yu et al. (2025). This field spans critical areas from climate change and emissions tracking IPCC (2023) to workplace safety and human rights Global Reporting Initiative (2023). Such knowledge is codified in technical frameworks and standards—major ones including GRI Initiative (2023), SASB Board (2023), TCFD Board (2017), ISSB International Sustainability Standards Board (2023), and CDP CDP Worldwide (2023)—all of which are constantly evolving.
@@
While LLMs show promise in processing complex ESG documents and providing relevant answers to user queries, their capabilities in this interdisciplinary domain remain largely unevaluated. Considering the high stakes involved, where incorrect responses about ESG requirements or metrics could lead to serious compliance violations or misguided sustainability initiatives, this assessment gap is particularly problematic. However, there is currently no comprehensive question-answering (QA) benchmark specifically designed to evaluate how well LLMs understand and reason about ESG concepts. Existing QA benchmarks either omit ESG topics entirely or address them only superficially. This gap leaves researchers and practitioners without a reliable way to measure and improve LLMs’ ESG knowledge comprehension and question-answering abilities.
@@
To close this gap, we present ESGenius, a curated benchmark that targets Multiple-Choice Questions (MCQs) answering as the core evaluation task for ESG and sustainability knowledge (Figure LABEL:fig:esgenius_pipeline). Our contributions are fourfold: (1) ESGenius benchmark: A comprehensive evaluation framework comprising two integrated components: (i) ESGenius-QA, a collection of 1,136 MCQs across various ESG pillars and sustainability topics, generated using state-of-the-art (SOTA) LLM approaches and validated by domain experts. Each question is explicitly mapped to supporting evidence from authoritative source texts, enabling transparent evaluation and facilitating RAG applications; and (ii) ESGenius-Corpus, a collection of 231 ESG documents and frameworks enabling efficient knowledge retrieval from 7 major authoritative sources. (2) Evaluation Protocol: We implement a comprehensive two-stage evaluation protocol consisting of zero-shot testing and RAG to systematically assess LLM capabilities. This assessment framework provides valuable insights into the current limitations and future potential of LLMs in understanding ESG. (3) Evaluation Analysis: Testing across a diverse set of 50 LLMs (0.5B to 671B) reveals significant performance gaps in zero-shot settings, with most models achieving 55–70% accuracy and the best model (o3) achieving a top score of 72.54%, as shown in Figure 2. However, models demonstrate substantial potential for improvement through RAG approaches, with DeepSeek-R1-Distill-Qwen-14B improving from 63.82% (zero-shot) to 80.46%. (4) Open Source Initiative: To foster community engagement and collaborative advancement, we have made our complete benchmark suite publicly available at https://github.com/ANGEL-NTU/ESGenius. This includes comprehensive documentation, evaluation code, model implementations, and the full ESGenius dataset. We also maintain an interactive web portal at https://angel-ntu.github.io/ESGenius featuring a real-time leaderboard and detailed performance visualizations through heatmap and one of its cells (Figures 9 and 10), enabling researchers to track progress and identify areas for improvement in ESG-focused language models.
@@
The remainder of this paper is organized as follows: §2 surveys prior efforts on ESG knowledge resources, question-answering benchmarks, and retrieval-augmented generation; §3 details the construction of the ESGenius benchmark, describing both the ESGenius-QA dataset and the ESGenius-Corpus; §4 outlines our experimental protocols—including zero-shot and RAG settings—and presents a comprehensive evaluation of 50 LLMs; §5 concludes the paper and highlights directions for subsequent research; §6 discusses the limitations of our benchmark, while §7 reflects on ethical considerations.
@@ 2 Related Work
Prior work at the intersection of LLMs and ESG has spanned diverse strands of research, from building domain-specific corpora and ontologies to designing evaluation benchmarks and retrieval-augmented generation methods. In this section, we review relevant literature across three key areas: (i) ESG and sustainability knowledge resources, which provide the taxonomies, corpora, and specialized models needed to capture domain-specific terminology; (ii) QA and evaluation benchmarks, which illustrate how standardized datasets have driven progress in measuring factual knowledge, reasoning, and domain expertise, while also exposing the absence of ESG-focused QA benchmarks; and (iii) RAG and knowledge grounding, which highlight how retrieval mechanisms enhance interpretability and factual accuracy in specialized domains such as finance and climate science.
@@ 2.1 ESG & Sustainability Knowledge Resources
ESG data have traditionally been guided by voluntary reporting frameworks such as the Global Reporting Initiative (GRI) and the Sustainability Accounting Standards Board (SASB). These frameworks define taxonomies of ESG topics (e.g., emissions, labor practices, board governance) and indicators, but the unstructured text in corporate sustainability reports poses significant challenges for computational use. To impose structure, researchers have proposed ontologies and knowledge bases for ESG. For instance, Zhou and Perzylo (2023) introduced OntoSustain, aligning GRI and EU standards (ESRS) to capture key sustainability indicators, though broad coverage and automated population remain difficult.
Several ESG-focused text corpora have emerged to support NLP research in this domain. One recent example is SusGen-3K, a dataset of 30k instances spanning multiple financial NLP tasks (e.g., sentiment, Q&A) and an ESG report generation task Wu et al. (2025). Meanwhile, Chang et al. (2024) built a benchmark of more than 8,000 labeled sentences from more than 14,000 corporate sustainability reports covering 36 ESG topics for topic and quality classification. Other datasets focus on more specific areas, such as the benchmark from Beck et al. (2025) for extracting greenhouse gas emissions, or on non-English languages, like the ESG-Kor dataset for information extraction in Korean Lee et al. (2024). These datasets remain relatively restricted, especially compared to general-domain corpora, and focus primarily on classification or generation tasks rather than MCQs question answering, thus motivating the need for broader coverage like ESGenius.
Prior NLP work on sustainability texts has largely centered around information extraction and text classification, evolving from early methods that relied on topic models or lexicon-based approaches Raghupathi et al. (2020) but often fell short in capturing nuanced ESG terminology. More recent efforts leverage LLMs, either by developing specialized models or by applying general models in sophisticated pipelines. In the former category, researchers have introduced domain-specific models like ClimateBERT, which continued BERT’s pre-training on 1.6M climate-related paragraphs for more effective risk disclosure classification Webersinke et al. (2021), and E-BERT, a set of models designed to optimize ESG reporting Zhang et al. (2025eb). In the latter category, approaches like ChatReport use GPT-based summarization and semantic search to assess a company’s report against TCFD recommendations Ni et al. (2023), while others have employed GPT-based pipelines to extract triples (company, action, ESG issue) and populate knowledge graphs from sustainability disclosures Bronzini et al. (2024). Collectively, these advanced approaches illustrate the promise of LLMs for ESG tasks but also highlight the need for standardized benchmarks for a thorough evaluation.
@@ 2.2 QA & Evaluation Benchmarks
Question answering (QA) benchmarks have driven many NLP advancements over the past decade, starting with the TREC QA tracks Voorhees et al. (1999) and continuing through datasets like SQuAD Rajpurkar et al. (2016). While these resources predominantly test general factual or reading comprehension, specialized QA datasets have emerged to measure higher-level reasoning and domain expertise. For example, OpenBookQA Mihaylov et al. (2018) challenged models to combine a small “open book” of scientific facts with common sense, while BoolQ Clark et al. (2019) featured yes/no queries requiring nuanced text understanding. Datasets such as DROP Dua et al. (2019) emphasize discrete reasoning skills, involving counting or arithmetic over paragraphs.
Further expansions have produced large-scale evaluation suites such as MMLU (Massive Multitask Language Understanding) Hendrycks et al. (2020) and MMMU (Massive Multi-Discipline Multimodal Understanding) Yue et al. (2024), which cover dozens of subjects at the college-exam level, along with domain-specific benchmarks like PubMedQA Jin et al. (2019) for biomedical research, JEC-QA Zhong et al. (2020) for legal question answering, and FinanceBench Islam et al. (2023) for finance. Yet despite these advances, there is still a notable lack of QA benchmarks dedicated to ESG and sustainability. To address this gap, ESGenius introduces a large collection of expert-validated MCQs spanning environmental, social, and governance topics. Drawing from authoritative frameworks (e.g., IPCC, GRI, SASB), ESGenius provides a high-stakes, real-world testbed tailored to ESG.
@@ 2.3 RAG and Knowledge Grounding
When tackling specialized or dynamically updated domains, RAG has emerged as a powerful method for accurate and interpretable QA. Lewis et al. (2020) introduced RAG, an approach that combines a parametric language model with a non-parametric memory of documents. This approach improved performance on knowledge-intensive QA tasks relative to purely parametric models, as it allows the LLM to ground responses in retrieved evidence.
Other retrieval-generation frameworks such as GraphRAG Edge et al. (2024), Fusion-in-Decoder Izacard and Grave (2021) and REALM Guu et al. (2020) illustrated diverse strategies for fetching relevant text chunks to guide the model’s reasoning. Recently, these methods have been applied to high-stakes sectors including finance and climate science. For instance, Vaghefi et al. (2023) integrated GPT-4 with a 3,000-page IPCC report to answer climate change questions more accurately. Similarly, Ni et al. (2023) used semantic search to retrieve TCFD-related segments from a company’s sustainability report, prompting an LLM to summarize potential climate risks. This grounding is critical in ESG contexts, where traceability and authoritative references are essential.
@@ 3 The ESGenius Benchmark
Developing reliable AI tools for the complex ESG and sustainability domain requires high-quality, domain-specific data. Our ESGenius benchmark provides a comprehensive foundation to evaluate and enhance LLMs in this context, drawing from globally recognized standards and authoritative sources. As illustrated in Figure LABEL:fig:esgenius_pipeline, the benchmark is built through a structured pipeline where ESG-related documents are digitized, LLMs generate candidate MCQs, and domain experts validate them to form ESGenius-QA, while the underlying source materials are systematically curated into the ESGenius-Corpus. Together, these two interconnected components ensure explicit linkages between questions and textual evidence, enabling transparent evaluation of ESG knowledge and supporting RAG methods. The detailed coverage of sources and statistics is presented in Appendix D.
@@ 3.1 ESGenius-QA
The ESGenius-QA dataset comprises 1,136 carefully curated MCQs with ground truth answers, validated by ESG domain experts. The questions systematically cover essential ESG topics across three core pillars including but not limited to: (i) Environmental: Climate change mitigation and adaptation, comprehensive carbon accounting (Scope 1, 2, 3 emissions), energy efficiency, water resource management, biodiversity conservation, waste reduction, and pollution control, etc; (ii) Social: Labor practices and standards, human rights protection, diversity and inclusion initiatives, workplace health and safety, community engagement and impact assessment, data privacy and protection, etc; (iii) Governance: Board structure and independence, executive compensation frameworks, business ethics and compliance, shareholder rights and engagement, enterprise risk management, and regulatory adherence, etc. The dataset is generated through our specialized pipeline (Appendix A.1) and undergoes rigorous validation by domain experts to ensure accuracy and relevance (§3.1).
@@
The ESGenius-QA dataset is structured to reflect the complexity and diversity of real-world ESG assessments through a standardized multiple-choice format. Each question contains four options (A-D) with a single correct answer, plus a dedicated “Not sure” option (Z) to capture model uncertainty. This design enables systematic evaluation of both performance and confidence. The dataset carefully balances conciseness with domain-specific precision. Questions and options employ precise ESG terminology while maintaining clarity and focus. Contextual prompts are provided where necessary for essential background. The specialized lexicon throughout authentically represents real-world ESG assessment scenarios. Table 1 presents a detailed token-level analysis of questions, options, and source text. Key terminology distributions are visualized through word clouds for questions (Figure 4), answer options (Figure 5) and the source text (Figure 6).
@@
Each question in ESGenius-QA is precisely mapped to relevant page(s) and text passages within the 7 authoritative source documents (GRI, SASB, IPCC, etc.) containing the information required for correct answers. ESG domain experts meticulously validate these mappings, ensuring both accuracy and relevance. This explicit linking between questions and their supporting evidence in source texts enables effective RAG approaches.
@@
To ensure the highest quality of questions and answers, we employ a multi-stage validation process (Appendix A.1). While Qwen-max-2025-01-25 was used in the initial phase to generate candidate QA pairs, the final benchmark was entirely curated by human experts: six independent reviewers with an average of five years’ experience in sustainability or NLP double-validated every question-answer pair, supported by three consulting industry experts with over ten years of relevant experience. Validators confirmed factual correctness against authoritative sources, clarity and unambiguity, plausibility yet incorrectness of distractors, and adequate context. Any question failing these criteria was refined or discarded, yielding a rejection rate of about 25% (1,136 of 1,519 MCQs retained). Importantly, Qwen-max-2025-01-25 was later evaluated only through its official API in a fully isolated session, ensuring no overlap or leakage between construction and evaluation. This expert-in-the-loop curation decisively eliminates circular dependency, and, consistent with standard practices in high-impact benchmarks, our method combines LLM-assisted generation with expert oversight to balance scale, quality, and domain relevance, while remaining open to validated alternatives to further reduce bias.